OpenStudy (anonymous):

If it requires 48.3 milliliters of 0.55 molar nitric acid to neutralize 15.0 milliliters of barium hydroxide, solve for the molarity of the barium hydroxide solution. Show all of the work used to solve this problem. Unbalanced equation: Ba(OH)2 + HNO3 yields Ba(NO3)2 + H2O

OpenStudy (anonymous):

@taramgrant0543664 @Photon336

OpenStudy (taramgrant0543664):

First step: You can find the moles of nitric acid using n=cv where n is the number of moles c is the concentration (aka molarity) so 0.55M v is volume and would need to be in litres not millilitres

OpenStudy (anonymous):

ok im here

OpenStudy (taramgrant0543664):

Second step: Balance chemical equation using the stoichiometric coefficients relate the HNO3 to the Ba(NO3)2 Using the ratio you can convert the moles of HNO3 to moles of Ba(NO3)2

OpenStudy (taramgrant0543664):

Third Step: After using the ratio you will have the moles of Ba(NO3)2 So you can use the formula c=n/v c is the concentration that you are trying to find n is the moles you found in the previous step v is the volume in litres so the 15mL but converted

OpenStudy (anonymous):

ok so the balanced equation is Ba(OH)2 + 2 HNO3 = Ba(NO3)2 + 2 H2O

OpenStudy (taramgrant0543664):

Yes it would be

OpenStudy (taramgrant0543664):

So the ratio is 2 to 1 For every 2 moles of HNO3 there is 1 mole of Ba(NO3)2 produced

OpenStudy (anonymous):

ok then what do i do

OpenStudy (taramgrant0543664):

SO since the ratio is 2:1 the moles you found in step 1 you divide that by 2 to get the moles of Ba(NO3)2 and then you can carry on to step 3

OpenStudy (anonymous):

C=N/V 0.55M=n/v

OpenStudy (anonymous):

i thought ml was volume

OpenStudy (taramgrant0543664):

mL is volume but since molarity is mol/L we have to convert the millilitres into litres 1L=1000mL

OpenStudy (anonymous):

but it's not 1000 ml

OpenStudy (taramgrant0543664):

That's the conversion 1mL=0.001L 48.3mL=? L

OpenStudy (anonymous):

0.0483L

OpenStudy (taramgrant0543664):

Yes using this and the 0.55M you can find n using the n=cv formula

OpenStudy (anonymous):

0.0483*0.55=0.26565

OpenStudy (taramgrant0543664):

Close I think you're missing a 0 in there

OpenStudy (anonymous):

0.026565

OpenStudy (photon336):

remember to add both volumes (V1 + V2) to get your final volume

OpenStudy (anonymous):

huh so 0.026565 isn't the answer?

OpenStudy (taramgrant0543664):

Yes so now we look at the ratio it was 2:1 so to get the moles of Ba(NO3)2 you divide the moles you just got by 2 to get the moles of Ba(NO3)2

OpenStudy (anonymous):

0.026565/2=0.0132825

OpenStudy (taramgrant0543664):

Yes so that would be the moles of Ba(NO3)2 So now you put that into the formula c=n/v and as @Photon336 had said you have to use the total volume so you have to add both together and convert that from millilitres to litres

OpenStudy (anonymous):

so 48.3+15.0?

OpenStudy (taramgrant0543664):

Yeo together that will be the final volume of the solution

OpenStudy (anonymous):

48.3+15.0=63.3 ---> 0.0633

OpenStudy (photon336):

yep

OpenStudy (anonymous):

so thats the final answer?

OpenStudy (photon336):

@lovingod786 they asked for concentration i think you need the number of moles too so it has to be moles/liter

OpenStudy (taramgrant0543664):

Nope you have to put it into the formula first c=n/v

OpenStudy (anonymous):

ok so c=0.26565/0.0633?

OpenStudy (taramgrant0543664):

0.0132825/0.0633=c

OpenStudy (anonymous):

0.21

OpenStudy (taramgrant0543664):

yep that would be correct

OpenStudy (photon336):

m\[M _{1}V_{1} = m _{2}V_{2} \] \[m_{2} = \frac{ M_{1}V{1} }{V_{2} }\] \[m_{2}] = moles of Ba(OH)2 \[V_{2}] = final volume.

OpenStudy (anonymous):

so thats the final answer then??

OpenStudy (anonymous):

??

OpenStudy (taramgrant0543664):

Yep

OpenStudy (anonymous):

OK THX

OpenStudy (anonymous):

You is a really good tutor or teacher or helper..you know whatever you clarify yourself as but you did a really good job!

OpenStudy (anonymous):

Well do not did

OpenStudy (taramgrant0543664):

Haha thanks! I try my best! I'm pretty bad at explaining things but this site has made me a lot better at explaining what I mean and @Photon336 is really good at explaining things too, probably one of the better people at explaining things in this section

OpenStudy (anonymous):

Yeah @Photon336 is good too when I first got here they use to help me

OpenStudy (photon336):

thanks guys